# Pathology-CoT: Learning Visual Chain-of-Thought Agent from Expert Whole Slide Image Diagnosis Behavior

### v1

@@ 4.5 Pathologist-o3 Agent Framework and Evaluation
    Reasoned Analysis and Summarization: Each candidate ROI is subsequently cropped at a higher magnification and processed by a central Vision-Language Model (VLM), termed the “Reasoning Module.” This module generates an explicit reasoning chain and detailed findings for each region. After analyzing all ROIs, the module synthesizes this information to produce a final diagnostic summary. Notably, unlike Pathology-CoT, pathologist-o3 does not require the VLM to rationalize the selection of an ROI, as this task is handled exclusively by the behavior predictor.
@@
Table S1: Performance Comparison using Different VLMs in Pathologist-o3 (Behavior Predictor Fixed). Best performed method(s) were highlighted in bold font.
@@
4.6 The Pathologist-o3 Agent: Behavior Predictor
To validate our dataset, we created an agent named Pathologist-o3. Its core navigational component is a task-specific Behavior Predictor, which is trained to propose diagnostically relevant ROIs on a new WSI. The Behavior Predictor is not a single, monolithic model. Instead, our framework envisions a library of predictors, each trained for a distinct, high-volume clinical task (e.g., melanoma screening, prostate cancer grading). For this paper, we focus on developing and evaluating one such predictor as a proof-of-concept: a model trained for the common and time-consuming task of identifying lymph node metastasis in colorectal cancer cases.
Our choice of a YOLO-based architecture for these individual Behavior Predictors is deliberate and grounded in clinical pragmatism. For this study, the predictor is a lightweight YOLOv5 [26] object detection model, chosen for its efficiency and ease of deployment. Its lightweight nature ensures that the model can be readily deployed in typical hospital IT environments, including on standard CPUs or within web-based viewers. This accessibility paves the way for future enhancements like on-the-fly incremental learning, where a predictor could be continuously personalized to a pathologist’s specific style.
@@ 4.6 The Pathologist-o3 Agent: Behavior Predictor
The Behavior Predictor was trained on the 1,521 ROI boxes extracted by the AI Session Recorder from 137 WSIs across all eight pathologists. It learns to identify areas with high diagnostic relevance based on expert behavior, which are not necessarily "abnormal" in a clinical sense. To ensure clinical validity and prevent data leakage, training was performed using a 5-fold cross-validation strategy, with splits made at the patient level.
@@
In the validation of the pathologist-o3 in Fig 2, Fig 4, Fig 5, prediction from 5 test fold is merged and used as Behavior Predictor’s output. In external validation on LN-CO2 dataset as Fig 3, fold-0 model is used to generate behavior bounding boxes prediction.
@@ 4.7.2 Comparison to State-of-the-Art Models
To benchmark our agent, we compared its diagnostic performance on the CRC lymph node task against a panel of leading VLMs (Fig. 2c). These baseline models were evaluated on their ability to perform the same task without the guidance of our Behavior Predictor. Full specifications of the models used and the dates of API access are detailed in Supplementary Table 1.
@@ 4.7.3 External Validation
To assess generalization, we performed a rigorous out-of-distribution test using the independent LN-CO2 cohort (Fig. 3). This dataset from different continental (Sweden) contains images from different scanners (Aperio and Hamamatsu) and at different magnifications (20x/40x), presenting a significant domain shift. The Behavior Predictor was trained exclusively on our single-center Stanford data and then evaluated end-to-end on the LN-CO2 dataset without any fine-tuning.
@@ 4.7.4 Behavioral Analyses
To deconstruct agent performance (Fig. 4), we conducted several analyses using the ground truth navigation path of a senior attending pathologist.
@@
  • Effect of Evidence Quantity: To simulate a chain-of-thought of varying length, we capped the maximum number of ROIs the agent could inspect at a given limit, $k$.
@@ 4.7.5 Practicality and Scalability Analysis
To assess practical viability (Fig. 5), we measured the per-slide cost and latency of our agent using two different VLMs: a high-performance model (Gemini 2.5 Pro) and a more cost-effective model (Gemini 2.5 Flash). Our cost calculations are based on public API pricing at the time of experimentation, which typically charges for both input and output tokens, with output tokens being more expensive. For image inputs, we standardized all ROIs to a 1024x1024 resolution to ensure consistent pricing per region, as many models charge based on pixel count. This analysis demonstrates that by carefully managing prompt length and standardizing image inputs, an agentic workflow can be made cost-effective, particularly when using smaller, faster models for less complex tasks. To analyze scalability, we compared three conditions across multiple VLM backbones: the baseline VLM without guidance, our agent guided by its learned behavior, and an oracle agent guided by the real behavior of an expert pathologist.
@@ Appendix A Appendix
Table S2: Overlap between VLM general capbility (MMMU) overall scores and Precision/Recall/Accuracy.
A.1 Supplementary Material: The Behavior Analyze Pipeline
In this section, we provide a detailed, step-by-step technical description of the Behavior Analyze pipeline used to preprocess raw pathologist interaction logs into a structured, VLM-friendly format. The entire pipeline is visually summarized in Figure S1, which illustrates the transformation of noisy, overlapping raw viewport data into a clean, normalized set of action bounding boxes. The implementation details described below correspond to the logic in the provided generate_via.py script.
@@ A.1 Supplementary Material: The Behavior Analyze Pipeline
A.1.1 Timing study design and fairness adjustment for Fig. 1c
We quantified the efficiency of the human-in-the-loop review relative to creating labels from scratch using a controlled timing study. Two users (User 1 and User 2) reviewed AI-drafted text for each region of interest (ROI) using the interface in Fig. S3. For every ROI, the system recorded wall-clock time from ROI load to decision. We summarize three per-round quantities: Verify (accepted with no edits), Revise (accepted after any edit), and their across-round Average. The experiment comprised 10 representative cases and 78 total conversational rounds.
To align with the chain-of-thought supervision used in the main task, users in the manual baselines were instructed to provide both a final diagnosis (e.g., lymph-node positive vs. negative) and a short rationale describing (i) abnormal/atypical features supporting the decision and (ii) normal structures present. Most AI drafts were accepted without edits; the observed revision rates were 15.59% for User 1 and 21.65% for User 2, consistent with over 80% of AI-generated text requiring no edits.
@@ A.1.2 Reviewer interface design
The reviewer tool (Fig. S3) was engineered to minimize edit latency while preserving an auditable record of decisions. The layout presents, from left to right, a slide thumbnail with the selected ROI box, a zoomed ROI view (and an optional cytology crop), and editable text panels for “Thumbnail Impression,” “Why Zoom,” and “Findings (ROI + Cyto).” The header shows case identifier, ROI index, progress, elapsed time, and the current decision state.
@@ A.1.3 Step 1: Initial Action Segmentation
For each identified action, a bounding box is created. For a StayInspect, the box is centered on the viewport. For a PanInspect, a minimal bounding box is calculated to encompass all viewports within the continuous panning sequence. The result is an initial list of VLMAction objects, as shown in the "Stay & Pan" column of Figure S1.
@@ A.1.4 Step 2: Filtering Overview Actions
Method: This step is performed by the filter_big_bboxes() function. Pathologists often zoom out to get their bearings, resulting in extremely large bounding boxes covering a substantial portion of the WSI. These actions are not useful for training a model to find specific ROIs. The function filters the action list by removing any VLMAction whose bounding box width is greater than a set fraction (e.g., 40% or 2/5) of the total WSI height (wsi_height). This effectively removes the "birds-eye view" actions, focusing the dataset on genuine inspection events. This is depicted in the "Remove Big" column of Figure S1.
@@ A.1.5 Step 3: Merging Overlapping Actions
Method: The merge_similar() function implements this logic. It iteratively compares all pairs of actions in the current list. If the Intersection over Union (IoU) between two action bounding boxes (calculate_iou()) exceeds a defined iou_threshold (e.g., 0.5-0.8), the two actions are considered part of the same inspection event. They are then merged into a single new VLMAction (merge_two_actions()). The new action’s bounding box is the union of the two original boxes, and its duration spans the full time from the start of the first action to the end of the second. This process repeats until no more merges are possible, resulting in a smaller set of more distinct actions, as seen in the "Merge" column of Figure S1.
@@ A.1.6 Step 4: Pruning Redundant Containing Actions
Method: This refinement is handled by the filter_mostly_contained_actions() function. It addresses cases where a pathologist inspects a large area (e.g., a lymph node) and then zooms in to inspect a specific feature within it (e.g., a cluster of tumor cells). In this scenario, the smaller, high-magnification action is more informative. The function compares every pair of actions. If a smaller action’s bounding box is mostly contained within a larger action’s box (i.e., the intersection area is >90% of the smaller box’s area, checked by check_containment_by_area()), the larger, containing action is marked as redundant and removed from the list. This prioritizes the most detailed and specific points of interest identified by the pathologist, as illustrated in the "Find Include" column of Figure S1.
@@ A.1.7 Step 5: Normalizing Action Bounding Boxes
Method: The final step, standardize_action_bboxes(), ensures that the input to the VLM is uniform. It maps each action’s bounding box to one of two standard sizes based on its original area, corresponding to a "5x" (larger box) or "10x" (smaller box) view. The function calculates the center of the final, pruned bounding box and resizes it to the standard dimension (e.g., wsi_height / 5.0 for a 5x view). This process also re-bins the action’s magnification label (magnification_bin) to match the standardized size. The output is a clean, normalized set of action bounding boxes of consistent sizes, centered on the pathologist’s key areas of focus. This final, structured output is shown in the "Normalize" column of Figure S1 and is ready for use in generating ground truth for model training.
A.2 Implementation Details on Dataset Construction
Seg-Zero Re-implementation. The evaluation of Seg-Zero [liu2025seg]’s 7B model is conducted using the official model checkpoint available in the public repository. Due to the unavailability of the 3B checkpoint, we re-train the model using the official codebase under the same experimental settings to ensure fair comparison. The prompts used for both models strictly adhere to the official implementation, as shown below. For consistency, the same prompt format is also adopted in our proposed model.
Original Prompt from Seg-Zero Please find “{Question}” with bbox and points. Compare the differences between objects and find the most closely matched one. Output the thinking process in <think> </think> and final answer in <answer> </answer> tags. Output the one bbox and points of two largest inscribed circles inside the interested object in JSON format. i.e., <think> thinking process here </think> <answer>“Bbox”: $[10,100,200,210]$, “Point 11”: $[30,110]$, “Point 22”: [35,180][35,180]</answer>
A.3 Ablation Study on Implementation
A.4 Prompts for VLM Analysis

### v2

@@ 4.5 Pathology-o3 Agent Framework and Evaluation
    Reasoned Analysis and Summarization: Each candidate ROI is subsequently cropped at a higher magnification and processed by a central Vision-Language Model (VLM), termed the “Reasoning Module.” This module generates an explicit reasoning chain and detailed findings for each region. After analyzing all ROIs, the module synthesizes this information to produce a final diagnostic summary. Notably, unlike Pathology-CoT, Pathology-o3 does not require the VLM to rationalize the selection of an ROI, as this task is handled exclusively by the behavior predictor.
@@
Table S1: Performance comparison using different VLMs in Pathology-o3 (Behavior Predictor Fixed). Best performed method(s) were highlighted in bold font.
@@
4.5.1 Statistic Analysis
To quantify the uncertainty of our performance metrics, we employed a bootstrap resampling method. The error bars presented for each model’s performance in Figure 2 and Figure 3 represent the 95% confidence intervals, which were derived from the 2.5th and 97.5th percentiles of 1,000 bootstrap iterations on the respective test sets.
To formally compare our Pathologist-o3 agent against the next-best performing models, we conducted a paired Student’s t-test on the distributions of performance differences obtained from the bootstrap procedure. Across both the internal validation set and the external LNCO2 cohort, Pathologist-o3 demonstrated a statistically significant improvement in overall performance over the second-best model, with p-values consistently below the significance threshold of 0.05 (p<0.05) for key metrics including accuracy, precision, and recall.
Table S2: Performance metrics of various models with 95% confidence intervals. Best performed method(s) were highlighted in bold font.
Model Prec. (%) Rec. (%) Acc. (%)
Gemini 33.3 (22.2 - 45.5) 18.2 (9.1 - 30.0) 60.0 (51.8 - 67.9)
o4-mini 55.0 (42.9 - 66.7) 45.8 (32.1 - 60.0) 65.6 (57.1 - 73.6)
GPT-4.1 42.3 (38.8 - 45.7) 91.7 (81.2 - 97.9) 50.0 (41.6 - 58.4)
QwenVL 37.3 (28.8 - 46.4) 52.1 (38.3 - 65.6) 49.2 (40.9 - 57.7)
Phi-4 41.0 (28.6 - 54.5) 33.3 (20.8 - 47.8) 57.0 (48.5 - 65.2)
Llama-4 75.0 (33.3 - 100.0) 6.2 (2.1 - 14.9) 64.1 (55.8 - 72.0)
InternVL 37.7 (34.5 - 41.0) 83.3 (70.8 - 93.8) 42.2 (34.0 - 50.4)
o3 46.7 (42.9 - 50.0) 87.5 (75.0 - 95.8) 57.8 (49.3 - 66.2)
Ours 84.5 (78.9 - 89.5) 100.0 75.4 (67.5 - 82.5)
Table S3: Performance metrics on the external LNCO2 cohort with 95% confidence intervals. Best performed method(s) were highlighted in bold font.
Model Acc. (%) Prec. (%) Rec. (%)
Gemini 62.0 (56.7 - 67.3) 23.5 (7.7 - 46.2) 7.8 (2.4 - 17.1)
Qwen 44.5 (39.3 - 50.2) 10.9 (5.3 - 18.4) 46.3 (31.7 - 61.0)
GPT-4o 40.3 (34.9 - 45.8) 9.8 (4.8 - 17.1) 51.5 (36.6 - 65.9)
o3 40.2 (34.9 - 45.5) 15.5 (9.3 - 23.3) 82.9 (70.7 - 92.7)
Pathologist-o3 69.4 (64.2 - 74.5) 62.9 (48.6 - 77.1) 97.6 (87.8 - 100.0)
4.6 The Pathology-o3 Agent: Behavior Predictor
To validate our dataset, we created an agent named Pathology-o3. Its core navigational component is a task-specific Behavior Predictor, which is trained to propose diagnostically relevant ROIs on a new WSI. The Behavior Predictor is not a single, monolithic model. Instead, our framework envisions a library of predictors, each trained for a distinct, high-volume clinical task (e.g., melanoma screening, prostate cancer grading). For this paper, we focus on developing and evaluating one such predictor as a proof-of-concept: a model trained for the common and time-consuming task of identifying lymph node metastasis in colorectal cancer cases.
Our choice of a YOLO-based architecture for these individual Behavior Predictors is deliberate and grounded in clinical pragmatism. For this study, the predictor is a lightweight YOLOv8 [29] object detection model, chosen for its efficiency and ease of deployment. Its lightweight nature ensures that the model can be readily deployed in typical hospital IT environments, including on standard CPUs or within web-based viewers. This accessibility paves the way for future enhancements like on-the-fly incremental learning, where a predictor could be continuously personalized to a pathologist’s specific style.
@@ 4.6 The Pathology-o3 Agent: Behavior Predictor
The Behavior Predictor was trained on the 1,521 ROI boxes extracted by the AI Session Recorder from 137 WSIs across all eight pathologists. It learns to identify areas with high diagnostic relevance based on expert behavior, which are not necessarily “abnormal” in a clinical sense. To ensure clinical validity and prevent data leakage, training was performed using a 5-fold cross-validation strategy, with splits made at the patient level.
@@
In the validation of the Pathology-o3 in Figure 2, Figure 4, Figure 5, prediction from 5 test fold is merged and used as Behavior Predictor’s output. In external validation on LNCO2 dataset (Figure 3), fold-0 model is used to generate behavior bounding boxes prediction.
To evaluate how effectively our model replicates the navigation behavior of expert pathologists, we define two primary metrics: Behavior Efficiency and Behavior Completeness. These metrics assess the model’s ability to predict the pathologist’s next area of focus, represented by bounding boxes corresponding to viewed regions. A predicted bounding box is considered a successful identification or “hit” if either its Intersection over Union with an expert-visited region surpasses a predefined threshold (0.3), or if one bounding box is contained within the other. This inclusive criterion is adopted to reasonably account for differing magnification preferences among pathologists, where a region of interest might be correctly identified but captured with a bounding box of a different scale, which could otherwise lead to a low IoU despite accurate localization.
Behavior Efficiency. This metric quantifies the precision of the model’s predicted areas of focus. It reflects how many of the model-proposed regions are indeed relevant to the expert’s examination path. High Behavior Efficiency indicates that the model minimizes suggestions for irrelevant areas, thus potentially leading to a more streamlined and efficient diagnostic workflow. It is calculated as:
$\text{Behavior Efficiency}=\frac{\text{Number of Hits}}{\text{Total Number of Model-Predicted Regions}}$
Behavior Completeness. This metric measures the recall of the model’s predictions. It assesses the proportion of expert-visited regions that the model successfully identifies. High Behavior Completeness suggests that the model comprehensively covers the areas deemed important by the expert pathologist, minimizing missed diagnostically relevant regions. It is calculated as:
[…]
@@ 4.7.2 Comparison to State-of-the-Art Models
To benchmark our agent, we compared its diagnostic performance on the CRC lymph node task against a panel of leading VLMs (Figure 2c). These baseline models were evaluated on their ability to perform the same task without the guidance of our Behavior Predictor. For reference we also provide their general vision capability in Table S4. Better performance on general vision task not results in more accurate diagnosis (Llama 4 vs Qwen). Full specifications of the models used and the dates of API access are detailed in Table S5.
@@ 4.7.3 External Validation
To assess generalization, we performed a rigorous out-of-distribution test using the independent LNCO2 cohort (Figure 3). This dataset from different continent (Sweden) contains images from different scanners (Leica Aperio and Hamamatsu) and at different magnifications (20x/40x), presenting a significant domain shift. (All slides from training cohort were scanned by Leica Aperio Scanner with 40x magnification at 0.25 µm per pixel). The Behavior Predictor was trained exclusively on our single-center Stanford data and then evaluated end-to-end on the LNCO2 dataset without any fine-tuning.
@@ 4.7.4 Behavioral Analyses
To deconstruct agent performance (Figure 4), we conducted several analyses using the ground truth navigation path of a senior attending pathologist.
@@
  • Effect of Evidence Quantity: To simulate a Chain-of-Thought of varying length, we capped the maximum number of ROIs the agent could inspect at a given limit.
@@ 4.7.5 Practicality and Scalability Analysis
To assess practical viability (Figure 5), we measured the per-slide cost and latency of our agent using two different VLMs: a high-performance model (Gemini 2.5 Pro) and a more cost-effective model (Gemini 2.5 Flash). Our cost calculations are based on public API pricing at the time of experimentation, which typically charges for both input and output tokens, with output tokens being more expensive. For image inputs, we standardized all ROIs to a 1024x1024 resolution to ensure consistent pricing per region, as many models charge based on pixel count. This analysis demonstrates that by carefully managing prompt length and standardizing image inputs, an agentic workflow can be made cost-effective, particularly when using smaller, faster models for less complex tasks. To analyze scalability, we compared three conditions across multiple VLM backbones: the baseline VLM without guidance, our agent guided by its learned behavior, and an oracle agent guided by the real behavior of an expert pathologist.
Author Contributions
S.W. and Z.H. conceived the study. S.W., R.W and Y. L. developed the methodology. S.W. and R.W. performed the experiments. S.W., J.S., and Z.H. curated and acquired the data. C.H., S.K., and J.S. provided expert pathology annotation and feedback. S.W. and Z.H. wrote the manuscript. Z.H. supervised the study. All authors discussed the results and approved the final manuscript.
Data Availability
Processed behavior data from 8 pathologists on the Stanford CRC lymph node dataset, including all intermediate bounding boxes, cropped ROIs, slide thumbnails, and the Chain-of-Thought dataset, will be made publicly available at https://github.com/zhihuanglab/Pathology-CoT. Whole-slide images and expert annotations from the LNCO2 dataset are available via the AIDA Data Hub at https://datahub.aida.scilifelab.se/10.23698/aida/lnco2 upon application and approval.
Code Availability
Source code is publicly available at https://github.com/zhihuanglab/Pathology-CoT.
Conflict of Interests
None declared.
@@ Appendix A Appendix
Table S4: Overlap between VLM general capbility (MMMU) overall scores and diagnosis Precision/Recall/Accuracy. General vision language capability have domain gap to pathology image diagnosis. Better performance on general vision task not results in more accurate diagnosis. MMMU results are from https://mmmu-benchmark.github.io
A.1 The AI Session Recorder Behavior Analyze Pipeline
In this section, we provide a detailed, step-by-step technical description of the Behavior Analyze pipeline used to preprocess raw pathologist interaction logs into a structured, VLM-friendly format. The entire pipeline is visually summarized in Figure S1, which illustrates the transformation of noisy, overlapping raw viewport data into a clean, normalized set of action bounding boxes. The implementation details described below correspond to the logic in the provided code.
@@ A.1 The AI Session Recorder Behavior Analyze Pipeline
Figure S4: Pathology-CoT Dataset Creation Pipeline. Left: Pathologist interaction data (viewports, navigation actions like zoom-in/panning, task context) is collected. Center: Generating the “Thinking Label”. Contextual information (task, current view, next action) is used to prompt a large language model (LLM, e.g., Gemini) to generate candidate reasoning. A pathologist then reviews, corrects, or approves this output to create the final “Thinking Label”. Right: The collected actions and generated thinking labels are merged with corresponding image patches (thumbnails, zoomed regions) and system prompts into a structured conversational format suitable for model training or in-context learning.
Figure S5: Sankey Diagram of Pathology-CoT Data Distribution The diagram illustrates how two ROIs, Inspect (broad-field scan) and Peek (high-magnification look), are applied to analyze various histological and cytological features during lymph node metastasis diagnosis. The flow size represents the relative proportion of diagnostic attention directed from a specific method to each feature.
A.1.1 Timing study design and fairness adjustment
We quantified the efficiency of the AI Session Recorder + human-in-the-loop review relative to creating labels from scratch using a controlled timing study. Two users (User 1 and User 2) reviewed AI Session Recorder-drafted text for each region of interest (ROI) using the interface in Figure S3. For every ROI, the system recorded wall-clock time from ROI load to decision. We summarize three per-round quantities: Verify (accepted with no edits), Revise (accepted after any edit), and their across-round Average. The experiment comprised 10 representative cases and 78 total conversational rounds.
To align with the chain-of-thought supervision used in the main task, users in the manual baselines were instructed to provide both a final diagnosis (e.g., lymph-node positive vs. negative) and a short rationale describing (i) abnormal/atypical features supporting the decision and (ii) normal structures present. Most AI Session Recorder drafts were accepted without edits; the observed revision rates were 15.59% for User 1 and 21.65% for User 2, consistent with over 80% of AI-generated text requiring no edits.
@@ A.1.2 Reviewer interface design
The reviewer tool (Figure. S3) was engineered to minimize edit latency while preserving an editable record of decisions. The layout presents, from left to right, a slide thumbnail with the selected ROI box, a zoomed ROI view (and an optional cytology crop), and editable text panels for “Thumbnail Impression,” “Why Zoom,” and “Findings (ROI + Cyto).” The header shows case identifier, ROI index, progress, elapsed time, and the current decision state.
@@ A.1.3 Step 1: Initial Action Segmentation
For each identified action, a bounding box is created. For a StayInspect, the box is centered on the viewport. For a PanInspect, a minimal bounding box is calculated to encompass all viewports within the continuous panning sequence. The result is an initial list of VLMAction objects, as shown in the “Stay & Pan” column of Figure S1.
@@ A.1.4 Step 2: Filtering Overview Actions
Method: This step is performed by the filter_big_bboxes() function. Pathologists often zoom out to get their bearings, resulting in extremely large bounding boxes covering a substantial portion of the WSI. These actions are not useful for training a model to find specific ROIs. The function filters the action list by removing any VLMAction whose bounding box width is greater than a set fraction (e.g., 40% or 2/5) of the total WSI height (wsi_height). This effectively removes the “birds-eye view” actions, focusing the dataset on genuine inspection events. This is depicted in the “Remove Big” column of Figure S1.
@@ A.1.5 Step 3: Merging Overlapping Actions
Method: The merge_similar() function implements this logic. It iteratively compares all pairs of actions in the current list. If the Intersection over Union (IoU) between two action bounding boxes (calculate_iou()) exceeds a defined iou_threshold (e.g., 0.5-0.8), the two actions are considered part of the same inspection event. They are then merged into a single new VLMAction (merge_two_actions()). The new action’s bounding box is the union of the two original boxes, and its duration spans the full time from the start of the first action to the end of the second. This process repeats until no more merges are possible, resulting in a smaller set of more distinct actions, as seen in the “Merge” column of Figure S1.
@@ A.1.6 Step 4: Pruning Redundant Containing Actions
Method: This refinement is handled by the filter_mostly_contained_actions() function. It addresses cases where a pathologist inspects a large area (e.g., a lymph node) and then zooms in to inspect a specific feature within it (e.g., a cluster of tumor cells). In this scenario, the smaller, high-magnification action is more informative. The function compares every pair of actions. If a smaller action’s bounding box is mostly contained within a larger action’s box (i.e., the intersection area is >90% of the smaller box’s area, checked by check_containment_by_area()), the larger, containing action is marked as redundant and removed from the list. This prioritizes the most detailed and specific points of interest identified by the pathologist, as illustrated in the “Find Include” column of Figure S1.
@@ A.1.7 Step 5: Normalizing Action Bounding Boxes
Method: The final step, standardize_action_bboxes(), ensures that the input to the VLM is uniform. It maps each action’s bounding box to one of two standard sizes based on its original area, corresponding to a “5x” (larger box) or “10x” (smaller box) view. The function calculates the center of the final, pruned bounding box and resizes it to the standard dimension (e.g., wsi_height / 5.0 for a 5x view). This process also re-bins the action’s magnification label (magnification_bin) to match the standardized size. The output is a clean, normalized set of action bounding boxes of consistent sizes, centered on the pathologist’s key areas of focus. This final, structured output is shown in the “Normalize” column of Figure S1 and is ready for use in generating ground truth for model training.
A.2 Prompts for VLM Analysis